# Automated Black-box Prompt Engineering for Personalized Text-to-Image Generation

### v2

@@ Abstract
Prompt engineering is an effective but labor-intensive way to control text-to-image (T2I) generative models. Its time-intensive nature and complexity have spurred the development of algorithms for automated prompt generation. However, these methods often struggle with transferability across T2I models, require white-box access to the underlying model, or produce non-intuitive prompts. In this work, we introduce PRISM, an algorithm that automatically produces human-interpretable and transferable prompts that can effectively generate desired concepts given only black-box access to T2I models. Inspired by large language model (LLM) jailbreaking, PRISM leverages the in-context learning ability of LLMs to iteratively refine the candidate prompt distribution built upon the reference images. Our experiments demonstrate the versatility and effectiveness of PRISM in generating accurate prompts for objects, styles, and images across multiple T2I models, including Stable Diffusion, DALL-E, and Midjourney.
@@ 3.1 Problem Statement
$y^{\star}\left(\{x_{i}\}_{i=1}^{M}\right)=\operatorname*{arg\,max}_{y\in{\mathcal{Y}}}\sum_{i=1}^{M}s(x_{i},y),$ (1)
@@ 3.2 Algorithm
Our method, Prompt Refinement and Iterative Sampling Mechanism  (PRISM), is an iterative process that repeats a prompt refinement subroutine for $K$ iterations in $N$ parallel streams, where $N\times K$ is a predetermined compute budget. At iteration $k$, the $n$-th stream of PRISM  randomly selects a reference image $x_{k,n}$ from $\{x_{i}\}_{i=1}^{M}$ and uses ${\mathbf{F}}$ to sample a candidate prompt $y_{k,n}$ from $p_{\theta_{\mathbf{F}}}(y\mid x_{k,n})$. Then it queries ${\mathbf{G}}$ to generate a single $\hat{x}_{k,n}$ from $y_{k,n}$ with $p_{\theta_{\mathbf{G}}}(x\mid y_{k,n})$ and evaluate the prompt with ${\mathbf{D}}$ to obtain an in-iteration score $s^{\prime}(x_{k,n},y_{k,n})=\mathbf{D}(x_{k,n},\hat{x}_{k,n})$. At the end of the iteration, we use the generated $y_{k,n}$ and its score to update $p_{\theta_{\mathbf{F}}}(y\mid x)$. After the entire process, we collect the subset of $\{y_{c}\}_{c=1}^{C}$ generated throughout this process that has the $C$-best in-iteration scores. Then we re-evaluate this subset with the total score $\sum_{i=1}^{M}s(x_{i},y_{c})$ and return the prompt with the best total score. If there is a tie, then we return the prompt with the highest log likelihood. The pseudocode and an illustration for this algorithm are outlined in Algorithm 1 and Figure 2 respectively.
The key difference between PRISM  and prior methods is that PRISM  updates the entire sampling distribution of prompts, whereas prior works (Gal et al., 2023; Wen et al., 2023; Mahajan et al., 2023) directly update the tokens of a single prompt or the embeddings of the prompt. We believe that maintaining the whole prompt distribution is beneficial as text-to-image generation is not a one-to-one operation, i.e. an image can be described by multiple different text prompts and the same text prompt can correspond to multiple differently generated images. Having access to the whole distribution allows the method to sample a more diverse range of prompts without starting from scratch and may also help the optimization escape potential local optima.
@@ 3.3 Designing and updating $\mathbf{F}$ and $p_{\theta_{\mathbf{F}}}$
In general, it is not obvious what the joint or the conditional distribution of all text and images is, so some form of approximation is unavoidable. In the context of image generation, a natural choice of the image-conditioned text distribution is an image captioning model. Traditional captioning models, however, fall short in controlled image generation for two primary reasons: (1) The level of detail necessary for generating specific images far exceeds what generic captioning models provide (Liang et al., 2023); (2) effective prompts for T2I models are often not grammatically correct sentences but rather collection of phrases that describe the details about the image, which generic captioning models are not trained to generate. For example, in Figure 5, the second reference image is generated by the prompt “A broken robot lays on the ground with plants growing over it, somber, HD, hyper realistic, intricate detail” with Stable Diffusion, but a caption for this image will not include components like “HD” or “hyper realistic”. As a result, instead of “a good description of an image”, we wish to directly model “possible prompts that are used to generate this image”.
@@ 3.4 Designing the judge model $\mathbf{D}$
We have a wider range of choices for the judge model as long as it provides a notion of similarity between a pair of images. A simple solution is to use pre-trained discriminative models such as CLIP (Radford et al., 2021), and measure the distance in their embedding spaces. While these models have seen various degrees of success, they also come with inherent limitations – the discriminative objective (e.g., contrastive loss) does not incentivize the model to attend to fine-grained details, an issue similar to the shortcomings of using captioning models to generate prompts (Liang et al., 2023). Moreover, in image generation, the criteria of success can be nuanced and difficult to quantify through traditional similarity metrics yet can be effortlessly described in human language. Lastly, the similarity we wish to measure may only involve some part of the visual features (e.g. color), and not all applications share the same notion of similarity. If we want to use pretrained discriminative models, then we need to find a different model for each task, which can be inefficient.
@@ 4.4 Ablation Study
4.5 Prompt Editing
@@ 5 Conclusion
Ethic Statement
Just as LLMs are suceptible to being jailbroken or adversarially manipulated by malicious actors (Zou et al., 2023), our method may also be vulnerable to malicious intent, potential bias, or limitations in the base models. Therefore, we will implement necessary safeguards upon the public release of our code and are committed to keep up with future advancements in improving the safety of our method.
@@ C.1 Additional Qualitative Results
In Figure 10, 11, 12, 13 and 14, we provide additional qualitative showcases for subject-driven personalized T2I generation, style-driven personalized T2I generation, direct image inversion and prompt editing. We also provide an example of the iteration and refinement process as a conversation between all three components in PRISM in Figure 15.
@@ C.2 Flexible Model Choices
C.3 Additional Applications
In this section, we demonstrate two additional applications of PRISM, prompt distillation and multi-concept generation.
PRISM is particularly well-suited for multi-concept generation due to the human readability of its generated prompts. This feature allows for easy identification and composition of different components within a scene, enabling intuitive control over multi-concept results. Unlike PEZ, which does not provide explicit control over which part of the prompt corresponds to specific aspects of the image (e.g., content or style), PRISM allows for much clearer and more direct manipulation, which we demonstrate in Figure 17
Unlike PEZ, which requires an additional optimization process to generate distilled prompts, PRISM leverages its highly interpretable prompts and the capabilities of LLMs to simplify prompts effectively. By using in-context learning and straightforward instructing a VLM (in this case a GPT-4o model) with the prompt “Here is a prompt to this image with a text-to-image model, make it more concise (less than $<$ token length constraint $>$ tokens) but keep all the descriptive details”, we achieve concise, distilled prompts without additional computational overhead, as we can observe in Figure 18.

### v3

@@ Abstract
Prompt engineering is an effective but labor-intensive way to control text-to-image (T2I) generative models. Its time-intensive nature and complexity have spurred the development of algorithms for automated prompt generation. However, these methods often struggle with transferability across T2I models, require white-box access to the underlying model, or produce non-intuitive prompts. In this work, we introduce PRISM, an algorithm that automatically produces human-interpretable and transferable prompts that can effectively generate desired concepts given only black-box access to T2I models. Inspired by large language model (LLM) jailbreaking, PRISM leverages the in-context learning ability of LLMs to iteratively refine the candidate prompt distribution built upon the reference images. Our experiments demonstrate the versatility and effectiveness of PRISM in generating accurate prompts for objects, styles, and images across multiple T2I models, including Stable Diffusion, DALL-E, and Midjourney. Our code is available via the project page here.
@@ 3.1 Problem Statement
$y^{\star}\left(\{x_{i}\}_{i=1}^{M}\right)=\operatorname*{arg\,max}_{y\in{\mathcal{Y}}}\sum_{i=1}^{M}s(x_{i},y)$ (1)
@@ 3.2 Algorithm
Our method, Prompt Refinement and Iterative Sampling Mechanism  (PRISM), is an iterative process that repeats a prompt refinement subroutine for $K$ iterations in $N$ parallel streams, where $N\times K$ is a predetermined compute budget. At iteration $k$, the $n$-th stream of PRISM  randomly selects a reference image $x_{k,n}$ from $\{x_{i}\}_{i=1}^{M}$ and uses ${\mathbf{F}}$ to sample a candidate prompt $y_{k,n}$ from $p_{\theta_{\mathbf{F}}}(y\mid x_{k,n})$. Then it queries ${\mathbf{G}}$ to generate a single $\hat{x}_{k,n}$ from $y_{k,n}$ with $p_{\theta_{\mathbf{G}}}(x\mid y_{k,n})$ and evaluate the prompt with ${\mathbf{D}}$ to obtain an in-iteration score $s^{\prime}(x_{k,n},y_{k,n})=\mathbf{D}(x_{k,n},\hat{x}_{k,n})$. At the end of the iteration, we use the generated $y_{k,n}$ and its score to update $p_{\theta_{\mathbf{F}}}(y\mid x)$. After the entire process, we collect the subset of $\{y_{c}\}_{c=1}^{C}$ generated throughout this process that has the $C$-best in-iteration scores. Then we re-evaluate this subset with the total score $\sum_{i=1}^{M}s(x_{i},y_{c})$ and return the prompt with the best total score. If there is a tie, then we return the prompt with the highest log likelihood. The pseudocode and an illustration are outlined in Algorithm 1 and Figure 2 respectively.
The key difference between PRISM  and prior methods is that PRISM  updates the entire sampling distribution of prompts, whereas prior works (Gal et al., 2023; Wen et al., 2023; Mahajan et al., 2023) directly update the tokens or the embeddings of a single prompt. We believe that maintaining the whole prompt distribution is beneficial as text-to-image generation is not a one-to-one operation, i.e. an image can be described by multiple different text prompts and the same text prompt can correspond to multiple differently generated images. Having access to the whole distribution allows the method to sample a more diverse range of prompts without starting from scratch and may also help the optimization escape potential local optima.
@@ 3.3 Designing and updating $\mathbf{F}$ and $p_{\theta_{\mathbf{F}}}$
In general, it is not obvious what the joint or the conditional distribution of all text and images is, so some form of approximation is unavoidable. In the context of image generation, a natural choice of the image-conditioned text distribution is an image captioning model. Traditional captioning models, however, fall short in controlled image generation for two primary reasons: (1) The level of detail necessary for generating specific images far exceeds what generic captioning models provide (Liang et al., 2023); (2) effective prompts for T2I models are often not grammatically correct sentences but rather collection of phrases that describe the details about the image, which generic captioning models are not trained to generate. For example, in Figure 5, the second reference image is generated by the prompt “A broken robot lays on the ground with plants growing over it, somber, HD, hyper realistic, intricate detail” with Stable Diffusion, but a caption for this image will not include components like “HD” or “hyper realistic”. Hence, instead of “a good description of an image”, we wish to directly model “possible prompts used to generate this image”.
@@ 3.4 Designing the judge model $\mathbf{D}$
We have a wider range of choices for $\mathbf{D}$ as long as it provides a notion of similarity between a pair of images. A simple solution is to use pre-trained discriminative models such as CLIP (Radford et al., 2021), and measure the distance in their embedding spaces. While these models have seen various degrees of success, they also come with inherent limitations – the discriminative objective (e.g., contrastive loss) does not incentivize the model to attend to fine-grained details, an issue similar to the shortcomings of using captioning models to generate prompts (Liang et al., 2023). Moreover, in image generation, the criteria of success can be nuanced and difficult to quantify through traditional similarity metrics yet can be effortlessly described in human language. Lastly, the similarity we wish to measure may only involve some part of the visual features (e.g. color), and not all applications share the same notion of similarity. If we want to use pretrained discriminative models, then we need to find a different model for each task, which can be inefficient.
@@ 4.4 Ablation Study
4.5 Additional Applications
Prompt Editing
Multi-Concept Generation
PRISM is particularly well-suited for multi-concept generation due to the human readability of its generated prompts. This feature allows for easy identification and composition of different components within a scene, enabling intuitive control over multi-concept results. Unlike PEZ, which does not provide explicit control over which part of the prompt corresponds to specific aspects of the image, PRISM allows for much clearer and more direct manipulation, which we demonstrate in Figure 10
Prompt Distillation
Unlike PEZ, which requires an additional optimization process to generate distilled prompts, PRISM leverages its highly interpretable prompts and VLMs to simplify prompts effectively. By directly instructing a VLM (in this case GPT-4o) with the prompt “Here is a prompt to this image with a text-to-image model, make it more concise (less than $<$length$>$ tokens) but keep all the descriptive details”, we achieve concise, distilled prompts without additional computational overhead, as shown in Figure 11.
@@ 5 Conclusion
Broader Impact Statement
Just as LLMs are susceptible to being jailbroken or adversarially manipulated by malicious actors (Zou et al., 2023), our method may also be vulnerable to malicious intent, potential bias, or limitations in the base models. Therefore, we will implement necessary safeguards upon the public release of our code and are committed to keep up with future advancements in improving the safety of our method.
@@ C.1 Additional Qualitative Results
In Figure 12, 13, 14, 15 and 16, we provide additional qualitative showcases for subject-driven personalized T2I generation, style-driven personalized T2I generation, direct image inversion and prompt editing. We also provide an example of the iteration and refinement process as a conversation between all three components in PRISM in Figure 18.